Halothane is an anesthetic that is 12.17% C, 0.51 % H,  40.48% Br, 17.96% Cl, and 28.87% F by mass. What is the compound's molar mass if each molecule contains exactly one hydrogen atom?
Answer #1

we have mass of each elements as:

C: 12.17 g

H: 0.51 g

Br: 40.48 g

Cl: 17.96 g

F: 28.87 g

Divide by molar mass to get number of moles of each:

C: 12.17/12.01 = 1.0133

H: 0.51/1.008 = 0.506

Br: 40.48/79.9 = 0.5066

Cl: 17.96/35.45 = 0.5066

F: 28.87/19.0 = 1.5195

Divide by smallest to get simplest whole number ratio:

C: 1.0133/0.506 = 2

H: 0.506/0.506 = 1

Br: 0.5066/0.506 = 1

Cl: 0.5066/0.506 = 1

F: 1.5195/0.506 = 3

So empirical formula is:C2HBrClF3

This contain 1 hydrogen atom.

So, same is the molecular formula

Molar mass of C2HBrClF3,

MM = 2*MM(C) + 1*MM(H) + 1*MM(Br) + 1*MM(Cl) + 3*MM(F)

= 2*12.01 + 1*1.008 + 1*79.9 + 1*35.45 + 3*19.0

= 197.378 g/mol

Answer: 197.4 g/mol